A) The Solubility Product Constant for lead hydroxide is 2.8×10-16.   If lead hydroxide is dissolved in...

A)

The Solubility Product Constant for lead hydroxide is 2.8×10-16.  

If lead hydroxide is dissolved in water you can say that the equilibrium concentrations of lead and hydroxide ions are: ...
...A. High
...B. Moderate
...C. Low

The solubility of lead hydroxide in water is:  ...
...A. High
...B. Moderate
...C. Low

B)he Solubility Product Constant for aluminum chloride is > 1.  

If aluminum chloride is dissolved in water you can say that the equilibrium concentrations of aluminum and chloride ions are: ...
...A. High
...B. Moderate
...C. Low

The solubility of aluminum chloride in water is:  ...
...A. High
...B. Moderate
Answer #1

A)

i) Equilibrium concentrations of lead and hydroxide ions are

C. Low

Solubility equilibrium of Pb(OH)2 is

Pb(OH)2(s) <--------> Pb2+(aq) + 2OH-(aq)

Ksp = [Pb2+] [OH-]2

If solubility of Pb(OH)2 is represented by S

at equilibrium ,

[Pb2+] = S

[OH-] = 2S

S(2S)2 = 2.8 ×10-16M3

4S3 = 2.8 ×10-16 M3

S3 = 7.0 × 10-17M3

S = 4.12 × 10-6M

Therefore, at equilibrium

[Pb2+] = 4.12 ×10-6M

[OH-] = 2× 4.12 ×10-6 = 8.24 ×10-6M

ii) Solubility of lead hydroxide in water is

C. Low

B)

i) Equilibrium concentrations of aluminium and chloride

A. High

ii) The solubility of aluminium chloride in water is

A. High